# LongViTU: Instruction Tuning for Long-Form Video Understanding

### v1

@@ Abstract
This paper introduce LongViTU, a large-scale (~121k QA pairs, ~900h videos), automatically generated dataset for long-form video understanding. We developed a systematic approach that organizes videos into a hierarchical tree structure and incorporates self-revision mechanisms to ensure high-quality QA pairs. Each QA pair in LongViTU features: 1) long-term context (average certificate length of 4.6 minutes); 2) rich knowledge and condensed reasoning (commonsense, causality, planning, etc.); and 3) explicit timestamp labels for relevant events. LongViTU also serves as a benchmark for instruction following in long-form and streaming video understanding. We evaluate the open-source state-of-the-art long video understanding model, LongVU, and the commercial model, Gemini-1.5-Pro, on our benchmark. They achieve GPT-4 scores of 49.9 and 52.3, respectively, underscoring the substantial challenge posed by our benchmark. Further supervised fine-tuning (SFT) on LongVU led to performance improvements of $12.0\%$ on our benchmark, $2.2\%$ on the in-distribution (ID) benchmark EgoSchema, $1.0\%$, $2.2\%$ and $1.2\%$ on the out-of-distribution (OOD) benchmarks VideoMME (Long), WorldQA and OpenEQA, respectively. These outcomes demonstrate LongViTU’s high data quality and robust OOD generalizability.
Figure 1: Illustration of LongViTU.  The top row shows an example video sequence, with red boxes highlighting key clues for the posed question and yellow boxes marking objects in key frames related to the answer. The middle row emphasizes the primary advantages of our proposed LongViTU over previous datasets, along with a Q&A sample; refer to section 1 for further details. The bottom row displays predictions from canonical open-source and commercial VLMs, evaluated by GPT-4 against ground truth using our novel predefined criteria.
@@ 1 Introduction
We present LongViTU, a novel dataset for large-scale, long-form video understanding (see Figure 1). Compared to existing Video Question-Answering (VQA) datasets, LongViTU provides greater naturalness and diversity. The primary advantages over prior works are summarized below, see subsubsection 2.1.1 for a clearer view.
@@
    Diverse Real World Scenarios.  Existing datasets suffer from limitations in domain coverage or scene diversity, like Env-QA [17] and OpenEQA [35] rely on videos from virtual environments, which inherently introduce a substantial domain gap. Others, including EgoVQA [14], EgoTaskQA [22], EgoSchema [36], WorldQA [67], MoVQA [64], MovieChat [45], and CinePile [41], etc., despite utilizing real world videos, primarily feature limited or homogeneous scenes. In contrast, LongViTU leverages the comprehensive Ego4D dataset [19], enabling VQA tasks to encompass a broader spectrum of real-world scenarios with enhanced diversity and realism.
@@
    Explicit Timestamp Labels.  Previous datasets, such as Otter [25], Video-ChatGPT [34], InternVideo [51], VideoChat [28], MVBench [29], LLaVA-Video [66], and MM-Ego [60], lack explicit timestamp labels for QA-related events. This omission means the precise start and end times for each QA are undefined, despite multiple QAs being present in a single video. To address this, our hierarchical pipeline organizes video content into a tree structure, enabling QA generation at various granularities with explicit timestamps for each event. LongViTU provides precise temporal annotations for all QA events, facilitating accurate identification of key moments in ultra-long videos. This approach shifts away from turn-based QA paradigms by LLMs & VLMs, enabling streaming QA for extended video content.
@@
    Long Certificate Length.  In short datasets, such as NextQA [54] and ActivityNet-QA [62], the average certificate length (introduced in EgoSchema [36], which we adhere to) is typically less than 10 seconds. While longer datasets like WorldQA [67] reach average certificate length under 60 seconds, and EgoSchema [36] stays below 100 seconds. LongViTU achieving an average certificate length of 276.8 seconds (~4.6 minutes). For further statistical details, please refer to Figure 3.
@@
    Fine-grained Categorization.  Most datasets focus primarily on basic question types related to spatial elements (such as objects, attributes, locations, and states, etc.), lacking a detailed and structured categorization. While recent datasets like EgoTaskQA [22] and OpenEQA [35] have introduced categorized questions, their classifications remain broad. In contrast, LongViTU emphasizes the spatiotemporal dimensions, offering a fine-grained categorization that captures the intricate spatiotemporal details within video contexts. This is organized into three primary categories, along with more granular subcategories, illustrated in Figure 3(b).
@@
    Open-ended Precise QA.  Unlike existing datasets with multiple-choice formats or irrelevant answers, LongViTU delivers open-ended precise QAs. By integrating a self-revision mechanism, LongViTU refines QAs to ensure alignment with video content and removing redundancies.
By preserving these primary advantages during dataset construction, the final dataset comprises ~121k high-quality QA pairs within ~900 hours of videos. To the best of our knowledge, LongViTU is the first publicly available long-form video question-answeringing dataset featuring explicit QA-related timestamp annotations. Our main contributions are summarized as follows:
@@
    We propose a novel automatic pipeline to generate high-quality video question-answeringing data, mitigating several limitations of existing datasets: diverse real world scenarios, explicit timestamp labels, long certificate length, fine-grained categorization, and open-ended precise QA.
  • $\bullet$
    With our pipeline, we curate LongViTU, a large-scale high-quality pre-training dataset and benchmark aimed at advancing instruction tuning for long-form and streaming video understanding.
  • $\bullet$
    We conducted extensive experiments demonstrating the benefits of LongViTU for canonical open-source and commercial VLMs and providing insightful analysis.
@@ 2 The LongViTU Dataset
We developed a hierarchical pipeline for processing long-form videos by structuring the content into a tree format. This approach effectively generates QA pairs with explicit timestamps and extended certificate lengths, while preserving fine-grained spatial details and temporal relevance across diverse scales.
2.1 Dataset Pipeline
2.1.1 Stage I: Hierarchical Video Tree Construction
Frame Level.  Commencing at the frame level, we leverage InternLM-XComposer2 [13] to perform multi-frame dense captioning (sampled at 1 fps) across annotated events in the Ego4D [19]. Descriptions are structured as $\langle d_{f},t_{s}^{f},t_{e}^{f}\rangle$, containing frame-level text description $d_{f}$ and accurate timestamps $t_{s}^{f}$ and $t_{e}^{f}$ derived from Ego4D’s temporal annotations for each event provided by human annotators.
@@ 2.1.1 Stage I: Hierarchical Video Tree Construction
Event Level.  To eliminate redundant frame level text, we apply GPT-4 [1] to refine both manually annotated events from Ego4D and automated dense captions, yielding concise event level descriptions $\langle d_{e},t_{s}^{e},t_{e}^{e}\rangle$ with accurate temporal boundaries $t_{s}^{e}$ and $t_{e}^{e}$.
@@
Table 1: Comparison with previous datasets.  The video sources for each dataset are listed under "Base", where "Collection" indicates that videos are derived from various resources. Furthermore, ✗ denotes multiple-choice answers, while ✓∗∗ indicates open-ended answers, LongViTU is the first large-scale dataset designed for long-form and streaming video understanding with explicit timestamp labels. The video durations and the number of QA pairs are approximate.
Dataset Base Scenario Open-ended Answer Fine-grained Categorization Explicit Timestamp Video Duration QAs
EgoVQA [14] IU Multiview [55] real world    ✗ 10h 600
Env-QA [17] AI2-THOR [24] virtual env       ✓∗∗ 130h 85.1K
EgoTaskQA [22] LEMMA [21] real world 15h 40K
MoVQA [64] Collection movie 50h 22K
MovieChat [45] Collection movie 160h 13K
CinePile [41] Collection movie 420h 303K
OpenEQA [35] ScanNet [11] HM3D [40] virtual env 3h 1.6K
WorldQA [67] PVSG [59] real world 10h 1K
EgoSchema [36] Ego4D [19] real world 250h 5K
LLaVA-Video [66] Collection real world 2000h 960K
MM-Ego [60] Ego4D [19] real world 3000h 7M
LongViTU (ours) Ego4D [19] real world 900h 121K
Segment Level.  GPT-4 further organizes these events into segments in the hierarchical video tree $\mathcal{T}_{\text{video}}$, merging related events into segments summarized at the segment level, represented as $\langle d_{s},t_{s}^{s},t_{e}^{s}\rangle$.
where $S$ denotes the total number of segments.
Video Tree Formulation.  Drawing on the pipeline above, we formalize the hierarchical tree structure for long-form video content as follows:
$\mathcal{T}_{\text{video}}=\langle R,\{\mathcal{S}^{i}\}_{i=1}^{S}\rangle$ (4)
where $\mathcal{T}_{\text{video}}$ represents the hierarchical tree, with $R$ as the root node and nodes down to frames. Figure 2 illustrates the whole architecture.
2.1.2 Stage II: Long-Form QA Generation
Sliding Window.  A sliding window operation on video subtrees generates QAs by balancing spatial details and temporal sequences. A five-segment window captures long-term temporal relevance and short-term spatial features, questions are derived from the first three segments, and answers from the subsequent two. The hyperparameters are set based on the duration distribution of the Ego4D dataset and can be adjusted for specific task requirements.
QA Generation.  We generate QA pairs from subtrees within the sliding window, formalized as:
$\mathcal{QA}_{\text{video}}=\big{\langle}Q,A,\mathcal{D}_{s}\big{\rangle}$ (5)
Here, $\mathcal{QA}_{\text{video}}$ denotes the set of generated QA pairs for the video subtree $\mathcal{T}_{\text{video}}$. The notation $\mathcal{D}_{s}$ represents the selected segments via the sliding window, where each segment $\mathcal{D}_{s}=\big{\langle}d_{s}^{i},\mathcal{D}_{e}\big{\rangle}$ $(s=1,\dots,S)$ consists of event-level descriptions $\mathcal{D}_{e}=\big{\langle}d_{e}^{j},\mathcal{D}_{f}\big{\rangle}$ $(e=1,\dots,E)$. Each event further includes frame-level descriptions $\mathcal{D}_{f}=\big{\langle}d_{f}^{k}\big{\rangle}$ $(f=1,\dots,F)$.
Self-Revision.  In this stage, GPT-4 rigorously reviews the generated question-answer pairs to verify their consistency with the video content. It identifies and corrects any inconsistencies or inaccuracies, eliminates irrelevant or unnecessary information, and emphasizes the core elements of the QA pairs to mitigate redundancy and overly simplistic responses. Additional details on the prompts and human evaluation procedures for this process are provided in Appendix C.
2.2 Characteristics and Statistics
Key Advantages.  Our two-stage methodology for constructing LongViTU aligns with the hierarchical structure of the video tree, offering several advantages that significantly enhance the data quality and utility:
  • $\bullet$
    Explicit Timestamp Labels: Each tree node is annotated with precise event timestamps, enhancing the reliability of temporal analysis.
  • $\bullet$
    Long Certificate Length: Comprehensive representation of event sequences ensures robust QA generation by covering longer temporal spans.
  • $\bullet$
    Fine-Grained Categorization: By leveraging specific subtrees, the methodology optimizes input length for large language models (LLMs), enabling effective processing of extended temporal events, preservation of spatial details, and support for QA generation across detailed categories.
Figure 2: Pipeline of LongViTU.  We adopt a hierarchical pipeline that organizes video content into a tree structure, with subtrees encapsulating information at different temporal scales. This framework facilitates the generation of QA pairs with explicit timestamps, ensuring adaptability to varying contextual lengths. By summarizing content across multiple temporal levels (frame level, event level, segment level), our approach overcomes the challenge of excessively long input length for LLMs from long-form video. This enables LLMs to generate distinct types of questions, resulting in a fine-grained categorization aligned with the video content. To further enhance quality, a self-revision step refines results by removing redundancy and irrelevant prior information. For more details, please refer to section 2.
Duration Distribution.  The LongViTU dataset contains 1,833 videos, divided into 1,533 for training, 200 for validation, and 100 for testing, totaling approximately 900 hours. The average video duration is 29.3 minutes, ranging from 3.5 to 120.7 minutes with a standard deviation of 17.5 minutes, following a long-tail distribution (Figure 3(a)). Each video is accompanied by an average of 66 QA pairs, QA durations range from 6 to 1800 seconds, averaging 276.8 seconds, also exhibiting a long-tail pattern. Events and segments have average durations of 8.5 and 82 seconds, respectively. The dataset includes 121k QA pairs: 101k for training, 14k for validation, and 6k for testing.
Category Distribution.  The sunburst diagram in Figure 3(b) illustrates the distribution of question-answer pairs across three main categories: Spatiotemporal Understanding (55%), further divided into Object (12.2%), Attribute (10.7%), Location (15.5%), and Action (16.6%); Episodic Reasoning (24.4%), including Transition (8.1%), Interaction (3.4%), Causality (5.4%), and Motivation (7.5%); and Commonsense Inference (20.6%), comprising Planning (5.4%), Risk (2.7%), Function (6.4%), and Affordance (4.6%). Detailed categorization and examples are provided in Appendix E.
Computational Cost. We utilized 8 NVIDIA A100 GPUs (80GB) for dataset construction and all experiments. Dense captioning of all videos required approximately 1 week, while the generation pipeline consumed an average of 10k tokens per QA-pair via the OpenAI API [37]. Training session durations ranged from a minimum of 3 hours to a maximum of 3 days.
3 Experiments
We evaluated mainstream VLMs on the LongViTU test set, with instruction tuning conducted using our training set. The results demonstrate that LongViTU poses unique challenges for modern VLMs, whether through frame-based or sampling-based approaches. SFT significantly improved performance on both In-Distribution (ID) and several canonical Out-of-Distribution (OOD) benchmarks, highlighting the strong generalization and robustness of LongViTU.
@@ 3 Experiments
3.1 Setup
Settings and Baselines.  We fine-tuned multiple vision-language models (VLMs) on the complete LongViTU training set and evaluated their performance on the test set to address challenges in long-form video understanding, all input data was standardized in alignment with LLaVA [33]. Frame-based models such as mPLUG-OWL [61] and Video-LLaVA [31] uniformly sampled 8 frames per video as input. In contrast, LLaMA-VID [30], LongVA-DPO [65], LongVU [44], and Gemini-1.5-Pro [42] sampled 1 frame per second, while VideoAgent [15] sampled 1 frame every 2 seconds. All models, each with a uniform parameter size of 7B, generated answers to video-related questions. Evaluation employed a multi-level scoring system, where GPT-4 assessed the textual alignment between generated answers and ground truth based on predefined criteria, scores were averaged across subcategories to ensure a fair comparison.
Metrics and Benchmarks.  Traditional caption metrics are inadequate for evaluating open-ended questions. To address this limitation, we developed a multi-level scoring criteria that leverages GPT-4’s near-human capabilities in text comprehension and alignment. This approach rigorously evaluates the alignment between predicted answers and ground truth, ensuring that key aspects of each question are addressed. Hallucinations or irrelevant responses are penalized, while concise and accurate answers are rewarded with higher scores. Detailed scoring criteria and prompts are provided in Appendix C. Furthermore, all models were fine-tuned on the LongViTU training set and evaluated on the ID benchmark EgoSchema [36] and the OOD benchmarks VideoMME [16], WorldQA [67], and OpenEQA [35], achieving superior performance over baseline models.
3.2 Quantitative Analysis on LongViTU
The quantitative results on LongViTU detailed in subsection 3.2 reveal the following critical insights:
Effective Fine-Tuning.  Our supervised fine-tuning (SFT) led the open-source model LongVU to achieve a top GPT-4 score of $55.9$, surpassing the commercial Gemini-1.5-Pro’s $52.3$ at a dense sampling rate of 1 fps. This convergence between top open-source and commercial models highlights the challenges posed by LongViTU in facilitating comprehensive understanding of long-form video content.
Dense Sampling Challenge.  At a sampling rate of 1 fps, LLaMA-VID underperforms compared to mainstream open-source video language models in both zero-shot and fine-tuned settings. This gap reflects limitations in existing dense sampling strategies for capturing essential features in long-form video understanding.
LLM Reliance Bias.  The robust performance in text-only blind QA suggests a potential bias in QA systems favoring textual predictions. This phenomenon highlights over-reliance on textual intermediaries, allowing direct inference of answers from questions, further discussion in Appendix A.

### v2

@@ Abstract
This paper introduces LongViTU, a large-scale (~121k QA pairs, ~900h videos), automatically generated dataset for long-form video understanding. We propose a systematic approach that organizes videos into a hierarchical tree structure for QA generation and incorporates self-revision mechanisms to ensure high-quality QA pairs. Each QA pair in LongViTU features: 1) long-term context (average certificate length of 4.6 minutes); 2) rich knowledge and condensed reasoning (commonsense, causality, planning, etc.). We also offer explicit timestamp annotations of relevant events for each QA pair. We have conducted extensive human studies on LongViTU, and the results prove the quality of our dataset. To better evaluate the challenges posed by LongViTU’s emphasis on long-term context and condensed reasoning, we manually curate a subset of LongViTU into a benchmark. Evaluations using a state-of-the-art open-source model (LongVU), a proprietary model (Gemini-1.5-Pro), and human annotators yield GPT-4 scores of 49.9, 52.3, and 81.0, respectively, underscoring the substantial difficulty presented by LongViTU questions. Performing supervised fine-tuning (SFT) of LongVU and LLaVA-Video on LongViTU data results in average performance gains of 2.5% and 3.7%, respectively, across a suite of long video understanding benchmarks (EgoSchema, VideoMME-Long, MLVU, LVBench).
Figure 1: Illustration of LongViTU.  The top row shows an example video sequence, with red boxes highlighting key clues for the posed question and yellow boxes marking objects in key frames related to the answer. The middle row emphasizes the primary advantages of our proposed LongViTU over previous datasets, along with QA examples; refer to section 1 for further details. The bottom row displays predictions from canonical open-source and proprietary VLMs, evaluated by GPT-4 against ground truth using our novel predefined criteria.
Table 1: Comparison with Existing Video QA Datasets.  The video sources for each dataset are listed under "Video Source", where "Collection" indicates that videos are derived from various sources. Furthermore, MC denotes multiple-choice answers, while OE∗∗ indicates open-ended answers, LongViTU is the first large-scale dataset designed for long-form video understanding with explicit timestamp annotations. The average certificate length, video durations and the number of QA pairs are approximate.
Dataset Video Source Scenario Answer Types Condensed Reasoning Explicit Timestamp Certificate Length Video Duration QAs
EgoVQA [14] IU Multiview [54] real world   MC 10s 10h 600
Env-QA [17] AI2-THOR [24] virtual env    OE∗∗ 10s 130h 85.1K
EgoTaskQA [22] LEMMA [21] real world OE 10s 15h 40K
MoVQA [62] Collection movie MC 230s 50h 22K
MovieChat [44] Collection movie OE 60s 160h 13K
CinePile [40] Collection movie MC 60s 420h 303K
OpenEQA [35] ScanNet [11] HM3D [39] virtual env OE 30s 3h 1.6K
WorldQA [65] PVSG [58] real world OE 60s 10h 1K
EgoSchema [36] Ego4D [19] real world  MC 100s 250h 5K
LLaVA-Video [64] Collection real world OE 30s 2000h 960K
MM-Ego [59] Ego4D [19] real world OE 10s 3000h 7M
LongViTU (ours) Ego4D [19] real world OE 276.8s 900h 121K
@@ 1 Introduction
We present LongViTU, a new dataset for large-scale, long-form video understanding (see Figure 1) that can be generated fully autonomously using LLM. Compared to existing Video Question-Answering (Video QA) datasets, LongViTU emphasizes long video context and condensed reasoning. The primary advantages are summarized below, further details can be found in Table 1.
@@
    Long Certificate Length.  Many Video QA datasets such as NextQA [53] and ActivityNet-QA [60] have a relatively short average certificate length (introduced by EgoSchema [36] to characterize the length of the video context required to answer the question), typically less than 10 seconds. While datasets with longer videos like WorldQA and EgoSchema [36] [65] can reach an average certificate length of 60~100 seconds, LongViTU attains an average certificate length of 276.8 seconds (~4.6 minutes) thanks to our proposed hierarchical tree-structured representations of videos in the data construction pipeline, highlighting its focus on long-context video understanding. For further statistical details, please refer to Figure 3.
@@
    Condensed Reasoning.  Existing QA datasets [22, 35] primarily focus on fundamental spatial elements (e.g., objects, attributes, locations, and states) but do not sufficiently emphasize reasoning. To facilitate questions that require condensed reasoning, we introduce a structured reasoning taxonomy that explicitly guides question generation by categorizing each question according to the specific type of reasoning it demands (e.g., affordance, planning, risk, functionality, causality). By prompting the LLM to align questions with these targeted reasoning categories, we significantly enhance the dataset’s capacity to assess condensed and complex reasoning capabilities.
@@
    Explicit Timestamp Annotations.  Previous datasets, such as Otter [25], Video-ChatGPT [34], InternVideo [50], VideoChat [28], MVBench [29], LLaVA-Video [64], and MM-Ego [59], lack explicit timestamp annotations for QA-related events. This omission means the precise start and end times for the relevant video context of each QA are undefined, despite multiple QAs being present in a single video. To address this, our tree-structured video representation enables QA generation at various granularities with explicit timestamps for each event. LongViTU provides precise temporal annotations for all QA events, providing this useful information for identifying key moments in ultra-long video understanding.
The final LongViTU dataset comprises ~121k high-quality QA pairs within ~900 hours of videos. To the best of our knowledge, LongViTU is the first publicly available, automatically generated long-form video question-answering dataset featuring condensed reasoning and explicit QA-related timestamp annotations. Our main contributions are summarized as follows:
@@
    We propose a novel automatic pipeline for generating high-quality video question-answering data, with a emphasis on long video context and condensed reasoning.
@@
    With our pipeline, we curate LongViTU, a large-scale, high-quality pre-training dataset, accompanied by a manually verified benchmark subset aimed at advancing long-form video understanding.
@@
    We conducted extensive experiments on a broad range of long video understanding benchmarks, demonstrating substantial benefits of LongViTU for canonical open-source VLMs, and providing insightful analyses.
Figure 2: The construction of LongViTU.  We design an automatic pipeline to generate QA pairs from Ego4D [19] videos while addressing the challenge of long-form video comprehension. To mitigate the context length limitation of LLMs and grasp spatial-temporal information in long video context, we employ a hierarchical tree structure. This structure first condenses dense captions to refine event descriptions and then segments content into sequential events with finer temporal granularity. Such representations facilitate the generation of QA pairs with explicit timestamps and varying contextual lengths, ensuring adaptability across different temporal levels (frame level, event level, segment level). Additionally, we design prompting strategies to generate high-quality questions that require condensed reasoning. Finally, a self-revision step refines results by removing redundancy and irrelevant prior information. For more details, please refer to section 2.
@@ 2 The LongViTU Dataset
Overview.  From a high level, we extract key video details, such as captions and human narration, and send them into an LLM to generate QA pairs. However, these details from long-form videos usually exceed the context limit of LLMs. Also, lengthy details make it difficult for LLM to grasp spatial-temporal information and generate favorable QA pairs (e.g., long certificate length). To address this, we employ a hierarchical tree structure: first condensing dense captions to refine event descriptions, then segmenting content into sequential events with finer temporal granularity (detailed in subsubsection 2.1.1). Based on the tree-based video representation, we design several prompting strategies to facilitate generating high-quality questions that require long video context and condensed reasoning (detailed in subsubsection 2.1.2). Finally, we conduct extensive analysis and quality assessment to ensure its quality (detailed in subsection 2.2 and subsection 2.3).
2.1 Dataset Construction Pipeline
2.1.1 Stage I: Tree-structured Video Representations
We design a hierarchical tree structure that encapsulates video semantics across multiple levels of granularity. This framework integrates frame-level, event-level, and segment-level, culminating in a structured tree $\mathcal{T}_{\text{video}}$. A visual illustration can be found in Figure 2.
Frame Level.  Commencing at the frame level, we leverage InternLM-XComposer2 [13] to perform multi-frame dense captioning (sampled at 1 fps) across annotated events in the Ego4D [19]. Descriptions are structured as $\langle d_{f},t_{s}^{f},t_{e}^{f}\rangle$, containing frame-level text description $d_{f}$ and explicit timestamps $t_{s}^{f}$ and $t_{e}^{f}$ derived from Ego4D’s temporal annotations for each event provided by human annotators.
@@ 2.1.1 Stage I: Tree-structured Video Representations
Event Level.  To eliminate redundant frame level text, an LLM is used to summarize both manually annotated events from Ego4D and the above frame-level automated dense captions, yielding concise event level descriptions $\langle d_{e},t_{s}^{e},t_{e}^{e}\rangle$ with explicit temporal boundaries $t_{s}^{e}$ and $t_{e}^{e}$.
@@
Segment Level.  The LLM further organizes these events into segments in the hierarchical video tree $\mathcal{T}_{\text{video}}$, merging related events into segments summarized at the segment level, represented as $\langle d_{s},t_{s}^{s},t_{e}^{s}\rangle$.
where $S$ denotes the total number of segments.
Video Tree.  Drawing on the pipeline above, we formalize the hierarchical tree structure for long-form video content as follows:
$\mathcal{T}_{\text{video}}=\langle R,\{\mathcal{S}^{i}\}_{i=1}^{S}\rangle$ (4)
where $\mathcal{T}_{\text{video}}$ represents the hierarchical tree, with $R$ as the root node and nodes down to frames.
2.1.2 Stage II: QA Generation using Video Trees
With the hierarchical video representations, we fetch a sub-tree (effectively a collection of consecutive segments) and send its corresponding video details (captions, narrations) to the LLM, then prompt it to generate the required questions based on the given video content, as detailed below.
Sliding Window and QA Generation.  A sliding window strategy over video sub-trees is employed to systematically generate QA pairs that balance spatial precision and temporal context. Specifically, each input window comprises five segments (a hyperparameter directly influencing the average certificate length), from which the LLM is prompted to generate QA pairs, formalized as:
$\mathcal{QA}_{\text{video}}=\big{\langle}Q,A,\mathcal{D}_{s}\big{\rangle},$ (5)
where $\mathcal{QA}_{\text{video}}$ denotes the set of generated QA pairs for the video sub-tree $\mathcal{T}_{\text{video}}$. The notation $\mathcal{D}_{s}$ represents the selected segments via the sliding window, where each segment $\mathcal{D}_{s}=\big{\langle}d_{s}^{i},\mathcal{D}_{e}\big{\rangle}$ $(s=1,\dots,S)$ consists of event-level descriptions $\mathcal{D}_{e}=\big{\langle}d_{e}^{j},\mathcal{D}_{f}\big{\rangle}$ $(e=1,\dots,E)$. Each event further includes frame-level descriptions $\mathcal{D}_{f}=\big{\langle}d_{f}^{k}\big{\rangle}$ $(f=1,\dots,F)$.
Ensuring Long Certificate Length.  Besides the adding aforementioned descriptions from input window of video segments, we further design the prompt to instructs the LLM to formulate questions based solely on information in the last two segments (Ask Content), while requiring answers exclusively found in the first three segments (Memory Content, see full prompt in  subsection C.2). This design explicitly ensures that generated questions necessitate referencing substantial preceding video context to guarante certificate lengths.
Facilitating Condensed Reasoning.  We’ve found that guiding the LLM to explicitly categorize its generated questions according to a structured reasoning taxonomy significantly improves its ability to produce questions that require condensed reasoning. Specifically, we design a hierarchical taxonomy (illustrated in Figure 3(b)) that classifies questions by the type of reasoning involved (e.g., affordance, planning, risk, functionality, causality, etc.). During question generation, we prompt the LLM to assign each generated question to the appropriate category within this taxonomy (detailed in subsection C.5). This structured categorization encourages the model to focus on generating questions that inherently demand more condensed and targeted reasoning.
Self-Revision.  After generating the initial QA pairs, the LLM undergoes rigorous self-revision to ensure consistency. This includes correcting inaccuracies, removing extraneous information, and refining responses to reduce redundancy and hallucinations, thereby improving dataset reliability. Additionally, the LLM conducts a pure-text evaluation (without visual input) to detect and discard QA pairs with excessive textual bias. The full prompt can be found in subsection C.3.
2.2 Characteristics and Statistics
Duration Distribution.  The LongViTU dataset contains 1,833 videos, divided into 1,533 for training, 200 for validation, and 100 for testing, totaling approximately 900 hours. The average video duration is 29.3 minutes, ranging from 3.5 to 120.7 minutes with a standard deviation of 17.5 minutes, following a long-tail distribution (Figure 3(a)). Each video is accompanied by an average of 66 QA pairs; QA durations range from 6 to 1800 seconds, averaging 276.8 seconds, also exhibiting a long-tail pattern. Events and segments have average durations of 8.5 and 82 seconds, respectively. The dataset includes 121k QA pairs: 101k for training, 14k for validation, and 600 human-reviewed samples for testing.
Category Distribution.  The sunburst diagram in Figure 3(b) illustrates the distribution of QA pairs across the aforementioned hierarchical taxonomy designed by us; detailed categorization and examples are provided in Appendix F.
2.3 Quality Assessment
We recruit a pool of raters to assess the quality of LongViTU data from two aspects: 1) evaluate the quality of LongViTU questions using a predefined rubric; 2) compare the quality of automatically generated LongViTU against VideoMME [16], a video QA benchmark that is annotated exclusively by humans. The results are shown in Figure 4.
Rubric-based assessment.  For the first aspect, the raters are guided to score the questions using the following rubric:
  • $\bullet$
    Good: The video segment contains rich and diverse events, and the generated QA pair exhibits strong spatio-temporal coherence with at least one event.
  • $\bullet$
    Fair: The video segment is relatively simple, and while the QA pair is well-aligned with an event, it primarily captures either spatial or temporal aspects rather than both.
  • $\bullet$
    Poor: The QA pair does not align well with the video content (e.g., it could be answered without reference to the video or is overly generic). However, it does not contain factual errors or hallucinations.
Due to resource constraints, only 100 randomly selected questions from LongViTU are evaluated. As suggested by the rubric, Good denotes excellent questions, while Fair questions remain highly valuable for long-video QA. Even Poor does not indicate incorrect QA but rather reflects weaker video alignment, which can still contribute to training more robust models. Overall, approximately 46% of the QA pairs exhibit strong spatio-temporal coherence, validating the quality of LongViTU.
Comparative Assessment.  We randomly sample 100 questions each from our autonomously generated LongViTU dataset and the manually annotated VideoMME dataset. Evaluators are asked to determine whether the quality of each question from LongViTU is significantly better, significantly worse, or comparable (Tie) to its counterpart in VideoMME. Results indicate that approximately 64% of questions from LongViTU achieve comparable quality to VideoMME, demonstrating that our fully autonomously generated dataset closely approaches human-annotated quality.
Figure 4: Quality Assessment.  Rubric-based assessment of LongViTU and comparative assessment against VideoMME [16].
3 Experiments
In our experiments, we aim to answer the following:
  1. 1.
    Do LongViTU questions indeed impose substantial challenges (long-video context and condensed reasoning) to today’s video understanding models?
  2. 2.
    To which degree does training on LongViTU enhance the capability of mainstream Vision-Language Models (VLMs) to understand long videos?
To answer the first question, we primarily perform zero-shot evaluation of mainstream video understanding models on the LongViTU test set. To answer the second question, we first perform supervised fine-tuning (SFT) of a collection of these models on LongViTU, then evaluate them on a collection of long-video understanding benchmarks (including LongViTU test set). To comprehensively evaluate this capability, we consider both In-Distribution (ID) and several canonical Out-of-Distribution (OOD) benchmarks. [^1] ID benchmarks share the same video source (Ego4D) as LongViTU, whereas OOD benchmarks utilize alternative sources. Detailed experimental settings are described in subsection 3.1, and results across various benchmarks are presented in subsection 3.2 and subsection 3.3.